OpenStudy (anonymous):

Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-33*x^-2*y^2/165*x^-1)^3

OpenStudy (mertsj):

\[(\frac{-33x ^{-2}y^2}{165x ^{-1}})^3=(\frac{-1y^2}{5x})^3=\frac{-y^6}{125x^3}\]

OpenStudy (anonymous):

thank you, appreciate you helping me.

OpenStudy (anonymous):

i got (-y^2/x^3/5)^3

OpenStudy (anonymous):

which one is correct?